# Think or Step-by-Step? UnZIPping the Black Box in Zero-Shot Prompts

### v1

@@ Abstract
Zero-shot prompting techniques have significantly improved the performance of Large Language Models (LLMs). However, we lack a clear understanding of why zero-shot prompts are so effective. For example, in the prompt “Let’s think step-by-step,” is “think” or “step-by-step” more crucial to its success? Existing interpretability methods, such as gradient-based and attention-based approaches, are computationally intensive and restricted to open-source models. We introduce the ZIP score (Zero-shot Importance of Perturbation score), a versatile metric applicable to both open and closed-source models, based on systematic input word perturbations. Our experiments across four recent LLMs, seven widely-used prompts, and several tasks, reveal interesting patterns in word importance. For instance, while both ‘step-by-step’ and ‘think’ show high ZIP scores, which one is more influential depends on the model and task. We validate our method using controlled experiments and compare our results with human judgments, finding that proprietary models align more closely with human intuition regarding word significance. These findings enhance our understanding of LLM behavior and contribute to developing more effective zero-shot prompts and improved model analysis. [^1] The complete codebase and documentation of language model interactions will be publicly available upon publication.
@@
Nikta Gohari Sadr, Sangmitra Madhusudan, and Ali Emami Brock University, St. Catharines, Canada {zu22of, sm20pd, aemami}@brocku.ca
@@ 1 Introduction
Prompting techniques have recently transformed natural language processing, significantly improving Large Language Model (LLM) performance across various tasks. From Chain-of-Thought (CoT) Wei et al. (2022) to Self-Consistency Wang et al. (2023b) and Least-to-Most prompting Zhou et al. (2022), these methods have expanded LLM capabilities. Zero-shot prompts, which guide models without task-specific training examples, have emerged as particularly effective and practical Kojima et al. (2022). However, despite their success, we lack a clear understanding of why these prompts are so effective, particularly at the level of individual words.
@@
Despite the success of such zero-shot prompting techniques, a critical gap in our understanding remains: What is it about these prompts that is so effective, particularly at the level of individual words? In the above prompt, which term has a more substantial influence on the model output: ‘think’ or ‘step-by-step’? Human intuition may suggest that certain components of a prompt are more influential than others; however, it frequently fails to offer a comprehensive explanation of the underlying mechanisms that generate LLM responses.
Current methods for interpreting the significance of input words in LLMs typically rely on gradients Wallace et al. (2019); Yin and Neubig (2022); Ferrando et al. (2023); Wu et al. (2023) or attention mechanisms Modarressi et al. (2022); Tenney et al. (2020); Ferrando et al. (2022). However, these methods are significantly constrained. They frequently require access to internal model parameters, rendering them inapplicable to closed-source models. Furthermore, interpretations based on gradients or attention may encounter reliability problems Wang et al. (2020); Jain and Wallace (2019); Ethayarajh and Jurafsky (2021); Zhao et al. (2024); Bastings and Filippova (2020), raising questions about their efficacy in accurately describing model behavior.
Our paper introduces the ZIP score (Zero-shot Importance of Perturbation score), a novel metric for assessing the importance of words in zero-shot instructional prompts. This metric addresses the limitations of current interpretability methods by:
  1. 1.
    Applying to both open and closed-source models without accessing internal parameters.
  2. 2.
    Avoiding reliability issues of gradient and attention-based approaches.
  3. 3.
    Revealing word-level importance through diverse and contextually relevant perturbations.
Figure 1 provides an overview of the ZIP score methodology. We validate our approach through extensive controlled experiments, statistical analyses, and comparisons with human judgments in seven influential zero-shot prompts, four LLMs, and several classification and translation tasks. Our findings reveal consistent patterns in word importance across models and tasks, offering new insights into LLM behavior and potential improvements in prompt engineering.
@@ 2 Related Work
Attention-Based Methods interpret feature importance by analyzing the weighted sum of intermediate representations in neural networks Modarressi et al. (2022); Tenney et al. (2020); Ferrando et al. (2022). Although intuitive, these approaches have several drawbacks. Studies have shown that attention weights do not always correspond to the most important features for model predictions Jain and Wallace (2019) and may not correlate well with other feature importance measures Ethayarajh and Jurafsky (2021). Moreover, attention weights often contain redundant and mixed information, which can lead to unreliable explanations Bastings and Filippova (2020). Like gradient-based methods, attention-based approaches also require access to model internals, limiting their applicability to open-source models only.
Perturbation-Based Methods, applicable to both open and closed-source models, assess importance by perturbing inputs and observing output changes. Notable approaches include LIME Ribeiro et al. (2016), which creates local linear approximations of model behavior, and other dataset instance perturbations Zafar and Khan (2019). Recent research has extended these methods to few-shot demonstrations Liu et al. (2023) and system prompts Hackmann et al. (2024); Yin et al. (2023).
@@ 3.1 Formalization of the ZIP score
Let $P$ represent the original prompt, where $P=w_{1},w_{2},\dots,w_{n}$ and $w_{i}$ denotes the $i$-th word. [^2] Throughout this paper, we use “word” to refer to any token resulting from our space-based tokenization method.
@@
For each word $w_{i}$ in $P$, we generate a set of perturbations $\{P_{i1},P_{i2},\dots,P_{ik}\}$, through context-aware synonym replacement and co-hyponym replacement using GPT-4333GPT-4 consistently outperformed traditional tools like NLTK Bird et al. (2009) in generating contextually appropriate alternatives in our preliminary tests., and word removal. Synonyms provide similar meanings, co-hyponyms introduce related concepts, and removal tests word necessity. Each $P_{ij}$ represents a modified version of $P$ where $w_{i}$ has been replaced or omitted.
@@
We posit that a word’s importance is reflected in the model’s sensitivity to any meaningful change in that word, whether a subtle shift in meaning or a complete removal (see §7 for analyses and demonstrations of this phenomenon).
@@
    The Universal Sentence Encoder Cer et al. (2018) to measure semantic similarity between $P$ and $P_{ij}$, keeping those with >30% similarity. [^4] This threshold was chosen empirically to balance maintaining context while allowing meaningful alterations.
@@
Across the seven instructional prompts, this filtering process resulted in an average of 9.04 perturbed sentences per word. Detailed prompts and examples for perturbation creation and evaluation are provided in Appendix Sections A.6 and A.7.
@@
Example instance (t): Is 2971 prime or composite?
@@
We define a disagreement function $d$ that quantifies the difference between the model predictions for each perturbed prompt ($P_{ij}$) and the original prompt ($P$). This function varies according to the type of task $T$. For an instance $t$ of classification tasks:
@@ 3.2 Identifying Significantly Important Words
The Wilcoxon rank-sum test evaluates whether two independent samples (i.e., the outputs from the original re-prompts and a perturbed prompt) come from the same distribution. Words yielding a p-value < 0.05 are considered ‘significantly important’. We use a temperature of 0.5 for all prompts to balance determinism and variability in model responses.
@@ 4 Controlled Pre-Experiment
Prompt Design and Methodologies: We constructed 20 validation prompts (Table 2) with varied syntax and structure, such as “Print the digits 123, “Type the letter X, and “Print carrot with no additional text”. Each prompt contains a key word that directly determines the correct output - any alteration to this word should change the model’s response. For example, in “Print the digits 123”, only perturbations to “123” should affect whether the model outputs “123” correctly. Table 11 in A.2 demonstrates that these prompts successfully ensure only the key word predominantly influences GPT-4o mini’s output.
@@ 5 Experimental Setup
We evaluated seven zero-shot instructional prompts across various datasets, as detailed in Table 1. These prompts were selected based on their prevalence in the recent literature and their effectiveness across different tasks. We used a simple space-based tokenization method, where tokens are defined as sequences of characters separated by whitespace. [^7] This approach preserves contractions and compound words (e.g., “Let’s” as one token), which are often semantically important in instructional prompts. Dataset-specific prompts are included in the Appendix A.6 (Table 20).
@@
Four recent LLMs were used in our experiments: GPT-4o mini, GPT-3.5-turbo, Llama-2-70B-chat, and Mixtral-8x7B-Instruct-v0.1 OpenAI (2023); Touvron et al. (2023); Jiang et al. (2024). Each model was configured with a temperature of 0.5 to balance between deterministic outputs and allowing for some variability. For a detailed breakdown of estimated token usage per instructional prompt required for ZIP Score computation across tasks, see Appendix A.8.
@@
We conducted a study with 15 diverse, English-proficient participants to compare human and model word importance preferences. For each prompt, participants selected up to three words they deemed most important. The evaluation form is shown in Appendix A.4 Fig. 17. We used the Jaccard Index to measure inter-annotator agreement across participants, yielding an average of 0.4714, indicating moderate agreement. Jaccard Index was chosen over Fleiss’ Kappa to accommodate multiple word selections per participant.
@@ 6.2 How do ZIP scores vary across models for the CoT prompt?
Table 3 presents ZIP score heatmaps for the Chain-of-Thought prompt “Let’s Think Step-by-Step” across four models using the Big Bench dataset. Our analysis reveals:
@@
Results on other datasets are provided in Appendix A.1 Table 5.
@@ 6.3 How does word importance vary across different prompts and task types?
  • ZIP scores inversely correlate with model performance. For GPT-4o mini, AQUA-RAT (accuracy: 68.26%, highest ZIP: 28.44) shows higher ZIP scores than GSM8K (91.59%, 5.83) and Big Bench (96.93%, 3.80). This suggests instructional prompts have greater impact on challenging tasks, while their influence diminishes on tasks where the model already excels.
@@ 6.4 How does the part-of-speech distribution of important words vary across models?
  • Noun dominance: Nouns (NN) consistently rank as most important across all models (47.4%–62.5% of significant words).
@@ 6.5 How well do model assessments of word importance align with human judgments?
Our analysis reveals that proprietary models (GPT-4o mini and GPT-3.5-turbo) show strong alignment with human word importance judgments. Open-source models (Mixtral and Llama 2) demonstrate less alignment, with Llama 2 showing the worst alignment to human judgments. Across all models and human evaluations, “step-by-step” consistently emerges as highly important, while “Let’s” is generally less emphasized. Appendix A.4 Figures 11 - 16 provide similar analyses for the remaining prompts used in our study.
@@ 7 Qualitative Analysis
We examined how subtle changes, particularly synonym substitutions, alter the responses of GPT-4o mini in classification and translation tasks.
In classification tasks, replacing “irrelevant” with “unrelated” in the prompt “Feel free to ignore irrelevant information” caused the model to overlook critical textual information, including numbers expressed in word form, leading to computational errors. For the prompt “Take a deep breath and work on this problem step-by-step", removing “step-by-step” surprisingly improved the model’s accuracy and adherence to given information. In translation tasks, replacing “refine” with “enhance” in “Please detect the error type, and refine the translation” degraded translation quality, showing how synonym substitution greatly and sometimes unintuitively affects task interpretation.
These examples demonstrate that seemingly equivalent words can produce significantly different model outputs. This observation supports our dual approach of using both synonym/co-hyponym substitution as well as word removal to reveal different aspects of how models interpret prompt wording. Details and heatmaps are provided in Appendix A.5 Tables 15 - 18.
@@ 8 Conclusion
Interdependencies Among Prompt Constituents: This study evaluates the influence of individual words within prompts through isolation-based perturbations. However, it does not account for the interdependencies and contextual relationships between words that could collectively influence the output of LLMs. Future research should consider methods that can analyze the joint effects of multiple words, providing deeper insights into the model’s processing mechanisms, and improving the reliability of perturbation-based interpretability methods.

### v2

@@ Abstract
Zero-shot prompting techniques have significantly improved the performance of Large Language Models (LLMs). However, we lack a clear understanding of why zero-shot prompts are so effective. For example, in the prompt “Let’s think step-by-step,” is think or step-by-step more crucial to its success? Existing interpretability methods, such as gradient-based and attention-based approaches, are computationally intensive and restricted to open-source models. We introduce the ZIP score (Zero-shot Importance of Perturbation score), a versatile metric applicable to both open and closed-source models, based on systematic input word perturbations. Our experiments across four flagship models, seven widely-used prompts, and several tasks reveal that: (1) word importance varies significantly by task type, with mathematical tasks prioritizing structured instruction words; (2) proprietary models align more closely with human judgments than open-source ones; (3) nouns consistently show the highest importance (47.4%–65.9%); and (4) ZIP scores inversely correlate with model performance (|r| > 0.9), indicating stronger prompt effects on challenging tasks. These findings enhance our understanding of LLM behavior and contribute to developing more effective zero-shot prompts and improved model analysis. [^1] The complete codebase and documentation of language model interactions will be publicly available upon publication.
@@
Nikta Gohari Sadr, Sangmitra Madhusudan, and Ali Emami Brock University, Saint Catharines, Canada {zu22of, sm20pd, aemami}@brocku.ca
@@ 1 Introduction
Prompting techniques have revolutionized how we use Large Language Models (LLMs), significantly improving their performance across various tasks. From Chain-of-Thought (CoT) Wei et al. (2022) to Self-Consistency Wang et al. (2023b) and Least-to-Most prompting Zhou et al. (2022), these methods have expanded LLM capabilities. Zero-shot prompts, which guide models without task-specific training examples, have emerged as particularly effective and practical, offering broad generalization and multi-task abilities Kojima et al. (2022). However, despite their success, we lack a clear understanding of why these prompts are so effective, particularly at the level of individual words.
@@
Despite the success of such zero-shot prompting techniques, a gap in our understanding remains: What is it about these prompts that is so effective, particularly at the level of individual words? In the above prompt, which term has more influence on the model output: ‘think’ or ‘step-by-step’?
Current methods for interpreting input word significance in LLMs typically rely on gradients Wallace et al. (2019); Yin and Neubig (2022); Ferrando et al. (2023); Wu et al. (2023) or attention mechanisms Modarressi et al. (2022); Tenney et al. (2020); Ferrando et al. (2022). However, these methods require access to internal model parameters and face reliability issues Wang et al. (2020); Jain and Wallace (2019); Ethayarajh and Jurafsky (2021); Zhao et al. (2024); Bastings and Filippova (2020).
Following established work Choudhary et al. (2022); Saleem et al. (2022); Feng et al. (2018), we define word importance as the degree to which a word’s presence, absence, or modification impacts model performance on a given task. This definition captures both semantic and structural importance through systematic perturbations: a word is considered important if replacing it with semantically similar alternatives (synonyms), related concepts (co-hyponyms), or removing it altogether significantly affects the model’s output.
Our paper introduces the ZIP score (Zero-shot Importance of Perturbation score), a novel metric for assessing word importance in zero-shot instructional prompts that works with both open and closed-source models. Figure 1 provides an overview of the methodology. Our approach is validated through controlled experiments, statistical analyses, and is compared with human judgment across seven zero-shot prompts, four LLMs, and several classification and translation tasks.
Our key findings reveal that: (1) Word importance varies significantly by task type (e.g., "step-by-step" for mathematical tasks vs. "think" for reasoning tasks); (2) Proprietary models like GPT-4o mini align more closely with human judgments of word importance than open-source models; (3) Nouns consistently show the highest importance (47.4%–65.9% of significant words) across all models; and (4) ZIP scores inversely correlate with model performance (|r| > 0.9), indicating that prompts have a greater impact on challenging tasks.
@@ 2 Related Work
Attention-Based Methods interpret feature importance by analyzing the weighted sum of intermediate representations in neural networks Modarressi et al. (2022); Tenney et al. (2020); Ferrando et al. (2022). While intuitive, these approaches have several drawbacks. For one, the attention weights do not always correspond to the most important features for model predictions Jain and Wallace (2019) and may not correlate well with other feature importance measures Ethayarajh and Jurafsky (2021). Moreover, attention weights often contain redundant and mixed information, leading to unreliable explanations Bastings and Filippova (2020). Like gradient-based methods, these approaches also require access to model internals, limiting their applicability to open-source models.
Perturbation-Based Methods can analyze any LLM regardless of architecture accessibility by measuring how outputs change when inputs are modified. As Choudhary et al. (2022) describes, “a word (token) or a collection of words (tokens) are modified or removed from the input samples, and a resulting change is measured.”
Notable approaches include LIME Ribeiro et al. (2016), which creates local linear approximations of model behavior, and other dataset instance perturbations Zafar and Khan (2019). Recent research has extended these methods to few-shot demonstrations Liu et al. (2023) and system prompts Hackmann et al. (2024); Yin et al. (2023).
@@ 3.1 Formalization of the ZIP score
Let $P$ represent the original prompt, where $P=w_{1},w_{2},\dots,w_{n}$ and $w_{i}$ denotes the $i$-th word. [^2] Throughout this paper, we use “word” to refer to any token resulting from our space-based tokenization method. This may include contractions, compound words, or multi-word expressions.
@@
For each word $w_{i}$ in $P$, we generate a set of perturbations $\{P_{i1},P_{i2},\dots,P_{ik}\}$, through context-aware synonym replacement and co-hyponym replacement using GPT-4333GPT-4 consistently outperformed traditional tools like NLTK Bird et al. (2009) in generating contextually appropriate alternatives in our preliminary tests., and word removal. Synonyms provide similar meanings, co-hyponyms [^4] A co-hyponym is a word that shares the same category or group with another word, but both are specific examples of a broader term (hypernym). introduce related concepts, and removal tests word necessity. Each $P_{ij}$ represents a modified version of $P$ where $w_{i}$ has been replaced or omitted.
@@
Aligned with existing literature on perturbation-based methods Choudhary et al. (2022); Ivanovs et al. (2021); Saleem et al. (2022), we posit that a word’s importance is indicated by the model’s performance shifts in response to any meaningful alteration of that word, whether through subtle meaningful shifts or its complete removal. (see §7 for analyses and demonstrations of this phenomenon).
@@
    The Universal Sentence Encoder Cer et al. (2018) to measure semantic similarity between $P$ and $P_{ij}$, keeping those with >30% similarity. [^5] This threshold was chosen to balance maintaining context while allowing meaningful alterations (see table 24 in A.7).
@@
We maintained cost-effectiveness in using large language models by optimizing token usage and the output format, reducing the tokens needed for perturbation generation and selection. Additionally, our primary filter, using Universal Sentence Encoder, discards unrelated perturbations, thereby minimizing unnecessary API calls for grammatical verification [^6] For the CoT prompt, the entire perturbation generation and grammatical verification process uses fewer than 2150 input/output tokens, effectively balancing cost and quality.. Across the seven instructional prompts, this filtering process resulted in an average of 9.04 perturbed sentences per word. Detailed prompts and examples for perturbation creation and evaluation are provided in Appendix Sections A.6 and A.7.
@@
Example ($t$): Is 29 prime or composite?
@@
We define a disagreement function $d$ that measures the difference between the model predictions for each perturbed prompt ($P_{ij}$) and the original prompt ($P$). This function varies according to the type of task $T$. For instance $t$ of classification tasks:
@@ 3.2 Identifying Significantly Important Words
The Wilcoxon rank-sum test evaluates whether two independent samples (i.e., the outputs from the original re-prompts and a perturbed prompt) come from the same distribution. Words yielding a p-value < 0.05 are considered ‘significantly important’.
@@ 4 Controlled Pre-Experiment
Prompt Design and Methodologies: We constructed 20 validation prompts (Table 2) with varied syntax and structure, such as “Print the digits 123, “Type the letter X, and “Print carrot with no additional text”. Each prompt contains a key word that directly determines the correct output - any alteration to this word should change the model’s response. For example, in “Print the digits 123”, only perturbations to “123” should affect whether the model outputs “123” correctly. Table 11 (A.2) confirms these prompts isolate the key word’s influence on GPT-4o mini’s output.
@@ 5 Experimental Setup
We evaluated seven zero-shot instructional prompts across various datasets, as detailed in Table 1. These prompts were selected based on their proven effectiveness and generalization capabilities in recent literature. We used a simple space-based tokenization method, where tokens are defined as sequences of characters separated by whitespace. [^9] This approach preserves contractions and compound words (e.g., “Let’s” as one token), which are often semantically important in instructional prompts. Dataset-specific prompts are included in the Appendix A.6 (Table 21).
@@
Four models were used in our experiments: GPT-4o mini, GPT-3.5-turbo, Llama-2-70B-chat, and Mixtral-8x7B-Instruct-v0.1 OpenAI (2023); Touvron et al. (2023); Jiang et al. (2024). Each model was configured with a temperature of 0.5 to balance between deterministic outputs and allowing for some variability. For token usage estimates per prompt, see Appendix A.8.
@@
We conducted a study with 15 diverse participants proficient in English to compare human and model interpretations of the instructional language. For each prompt, participants selected up to three words that they deemed most important to solve a certain task. The evaluation form is shown in Appendix A.4 Fig. 17. We used the Jaccard Index to measure interannotator agreement across participants, yielding an average of 0.4714, indicating moderate agreement. [^10] Jaccard Index was chosen over Fleiss’ Kappa to accommodate multiple word selections per participant.
@@ 6.2 How do ZIP scores vary across models for the CoT prompt?
Table 3 presents ZIP score heatmaps for the CoT prompt “Let’s Think Step-by-Step” across models on Big Bench. Our analysis reveals:
@@
Results on all datasets are provided in Appendix A.1 Table 5.
@@ 6.3 How does word importance vary across different prompts and task types?
  • ZIP scores strongly correlate inversely with model performance (|r| > 0.9) across all prompts. For example, using CoT, GPT-4o mini shows higher ZIP scores on AQUA-RAT (accuracy: 68.26%, ZIP: 28.44) compared to GSM8K (91.59%, 5.83) and Big Bench (96.93%, 3.80) indicating that instructional prompts have greater impact on challenging tasks (Table 15).
@@ 6.4 How does the part-of-speech distribution of important words vary across models?
  • Noun dominance: Nouns (NN) consistently rank as most important across all models (47.4%–65.9% of significant words).
@@ 6.5 How well do model assessments of word importance align with human judgments?
Proprietary models (GPT-4o mini and GPT-3.5-turbo) align strongly with human judgments, while open-source models show less alignment, particularly Llama 2. “Step-by-step” emerges as highly important across all evaluations, while “Let’s” is consistently less emphasized. Similar analyses for remaining prompts are provided in Appendix A.4 Figures 11 - 16.
@@ 7 Qualitative Analysis
We examined how subtle changes alter GPT-4o mini’s responses across tasks. In classification tasks, replacing “irrelevant” with “unrelated” in “Feel free to ignore irrelevant information” caused the model to overlook critical textual information, including numbers expressed in word form, leading to computational errors. For “Take a deep breath and work on this problem step-by-step", removing “step-by-step” surprisingly improved the model’s accuracy. In translation tasks, replacing “refine” with “enhance” in “Please detect the error type, and refine the translation” degraded translation quality, showing how synonym substitution can significantly affect task performance.
These examples demonstrate that seemingly equivalent words can produce significantly different model outputs. This observation supports our dual approach of using both synonym/co-hyponym substitution as well as word removal to reveal different aspects of how models interpret prompt wording. Details and heatmaps are provided in Appendix A.5 Tables 16 - 19.
@@ 8 Conclusion
Interdependencies Among Prompt Constituents: This study evaluates the influence of individual words within prompts through isolation-based, a necessary simplification grounded in established practices to assess the independent significance of each word. However, it does not account for the interdependencies and contextual relationships between words that could collectively influence the output of LLMs. Future research should consider methods that can analyze the joint effects of multiple words, providing deeper insights into the model’s processing mechanisms, and improving the reliability of perturbation-based interpretability methods.
@@ A.3 ZIP Scores
Prompt Type Correlation (r) Significance (p)
0-CoT -0.9948 0.064
0-CoTB -0.9954 0.060
0-CoTR -0.9982 0.037*
0-IRR -0.9999 0.005**
0-PS -0.9970 0.049*
Table 15: Pearson correlations between ZIP scores and GPT-4’s accuracy across classification tasks for different zero-shot prompts. All correlations show strong negative relationships (r < -0.99), with several reaching statistical significance (*p < 0.05, **p < 0.01). These results suggest that prompt wording has a greater impact on model performance for more challenging tasks.
@@ A.7 Perturbations
Original Word Generated Candidates 20% 30% 40% 50%
Let’s We should (49.13%) Accept Accept Accept Reject
It is recommended that we (21.29%) Accept Reject Reject Reject
We can (51.57%) Accept Accept Accept Accept
First Before anything else (47.05%) Accept Accept Accept Reject
Right off the bat (25.66%) Accept Reject Reject Reject
Understand Perceive (58.85%) Accept Accept Accept Accept
Apprehend (42.68%) Accept Accept Accept Reject
Problem Dilemma (72.35%) Accept Accept Accept Accept
Hurdle (35.48%) Accept Accept Reject Reject
Difficulty (60.58%) Accept Accept Accept Accept
Devise Design (37.65%) Accept Accept Reject Reject
Draft (36.57%) Accept Accept Reject Reject
Set up (36.85%) Accept Accept Reject Reject
Plan Procedure (44.23%) Accept Accept Accept Reject
Strategy (57.06%) Accept Accept Accept Accept
Solve Crack (29.82%) Accept Reject Reject Reject
Tackle (40.75%) Accept Accept Accept Reject
Step-by-step Progressively (38.88%) Accept Accept Reject Reject
Phase by phase (30.38%) Accept Accept Reject Reject
Inch by inch (21.12%) Accept Reject Reject Reject
Accuracy 85% 95% 65% 40%
Table 24: Evaluation of semantic similarity thresholds (20%-50%) on the 0-PS prompt using 20 manually validated word variants. Each row shows an original word, its generated alternatives with similarity scores, and acceptance decisions across thresholds. Green indicates correct decisions (accept/reject), red indicates incorrect ones. The accuracy row demonstrates that 30% provides optimal performance (95%) in identifying meaningfully related perturbations.